During its first year of operations, the McCormick Company incurred the following manufacturing costs: Direct materials, $5 per unit, Direct labor, $2 per unit, Variable overhead, $4 per unit, and Fixed overhead, $324,000. The company produced 36,000 units, and sold 28,500 units, leaving 7,500 units in inventory at year-end. Income calculated under variable costing is determined to be $400,000. How much income is reported under absorption costing?

Multiple Choice

  • $400,000

  • $332,500

  • $724,000

  • $467,500
Answer #1

Answer

--Correct Answer = Option #4: $ 467,500

--Units sold is LESS than Units produced, and hence Income under Variable costing must be LESS than Absorption Costing income.

--Ending units cost of fixed overhead is to be added to Variable costing income statement.

--Fixed Overhead per unit = $ 324000 / 36000 units produced = $ 9 per unit

--Ending units = 36000 - 28500 = 7500 units

--Income under absorption costing = $ 400000 + (7500 units x $ 9)
= 400000 + 67500
= $ 467,500